OpenStudy (anonymous):

Is there a shortcut when foiling this equation: (6t^2-18t+12)^2

OpenStudy (anonymous):

I know that there is a shortcut when there are two terms within the bracket but I wonder if there is one for one with 3 terms

OpenStudy (anonymous):

Im honestly not sure... Sorry

OpenStudy (anonymous):

That was a profound answer so u deserve a medal

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Strobe u shld know the answer

OpenStudy (anonymous):

Other than noting each time you multiply unlike terms you do this twice, so can just multiply by 2 the first time, I wouldn't say there's many optimisations you can make. These can be done so quickly in your head there's no point grinding over finding the quickest way, though.

OpenStudy (mathteacher1729):

Factor out the six, and realize you're really factoring out a six SQUARED. That makes the inside three terms a bit easier to square. \[(6t^2-18t+12)^2 = \left(6(t^2-3t+2)\right)^2= 6^2 (t^2-3t+2)^2 = 6^2\left((t-1)(t-2)\right)^2\] And if you wanna go nuts, that last step equals... \[6^2(t-1)^2\cdot(t-2)^2\] So yeah. Not exactly a walk through a daisy field, but slightly more manageable. :)

OpenStudy (anonymous):

ya thanks everyone

OpenStudy (anonymous):

HEY

hero (hero):

I did it

OpenStudy (anonymous):

what did u do?

hero (hero):

I'm applying to a new school. I had a discussion with the admissions counselors and I decided to go through with it

OpenStudy (anonymous):

WHAT????

OpenStudy (anonymous):

U r so fast

OpenStudy (anonymous):

Where r u applying to?

hero (hero):

I don't mess around

hero (hero):

Ottawa University

OpenStudy (anonymous):

but the new term only starts like in a month and a half

hero (hero):

I know. Now is a good time to start preparing

OpenStudy (anonymous):

oh so next term you are gonna sign up with ottawa and this term you r gonna be where u r

hero (hero):

Yes

OpenStudy (anonymous):

But I thought they don't offer wht u want?

hero (hero):

Actually, the next term begins 7 Jan 2012

OpenStudy (anonymous):

So what courses r u gonna take?

hero (hero):

I can just work on getting a degree in Math for right now.

hero (hero):

The pace is appropriate for me. 8 weeks

hero (hero):

I can get done with courses quick and focus the extra time doing other things I need to do.

OpenStudy (anonymous):

but what courses are u taking?

hero (hero):

Math Science courses right now, but each course is 3 and 1/2 weeks long and too dense.

OpenStudy (anonymous):

no what courses r u gonna take next term?

hero (hero):

Whatever courses are available. If you want, we can take the same courses.

OpenStudy (anonymous):

but i am still in the begginning

OpenStudy (anonymous):

I am taking next term linear algebra and calc 2

hero (hero):

Okay, I've never taken linear algebra before but I can study up on it before I take it

hero (hero):

I can take calc II again because I need to brush up

OpenStudy (anonymous):

that wld be a waste of time and money

OpenStudy (anonymous):

YOu shld get in touch with them

OpenStudy (anonymous):

they may make u take like some kind of english course beforehand

OpenStudy (anonymous):

They made me take this course first to get used to an online course

hero (hero):

No they won't. I already have experience taking online courses. It's on my transcript

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so get in touch with them

hero (hero):

and I've already taken English courses at this university

hero (hero):

I already spoke with them already

OpenStudy (anonymous):

idk u gotta ask them

OpenStudy (anonymous):

u did????

hero (hero):

Yes, I told you that already

OpenStudy (anonymous):

oh so what did they say that u r in?

hero (hero):

All I have to do is apply and send them my transcripts

OpenStudy (anonymous):

oh so u r gonna do that. Are u totally doing it online?

hero (hero):

yes

hero (hero):

I'll be doing the same thing you're doing

OpenStudy (anonymous):

oh cool

hero (hero):

I can't believe you've been on here all day

OpenStudy (anonymous):

kind of. I went away for like 3 hrs but now I am back and don't know how to foil this. KIND OF EMABARRASSING

hero (hero):

I can show you how to do it

hero (hero):

I can show you how to do almost anything

OpenStudy (anonymous):

well my lil sis is showing me now

hero (hero):

Oh, okay :D

hero (hero):

Another way to do it in case you're interested